OpenStudy (anonymous):

Tony is training to run a marathon. He has already run 2 miles today and is running additional miles at 6 miles per hour. Write an equation that represents the miles, M, that Tony runs, where h represents the number of hours.

OpenStudy (anonymous):

A. M = 6h + 2 B. M = 6h – 2 C. M = 2h + 6 D. M = –2h + 6

OpenStudy (texaschic101):

m= 2 + 6h

OpenStudy (texaschic101):

oops..or m = 6h + 2

OpenStudy (anonymous):

Thank you very much! Would you like a medal?

OpenStudy (texaschic101):

thats up to you :)